OpenStudy (antonio_xx2):

Tori's twelve-ounce glass of soymilk has 9 grams of protein. Brady's sixteen-ounce glass has twelve grams of protein. What is the ratio of the protein in Brady's drink to the protein in Tori's drink? one colon one three colon four four colon three sixteen colon four explaine please

OpenStudy (antonio_xx2):

@kirbykirby

OpenStudy (kirbykirby):

Tori: 12 oz has 9 grams Brady: 16 oz has 12 g What happens if Tori's volume increase to 16 (while proportionately increasing the grams of protein) ? \[\frac{12}{9}=\frac{16}{x}\\ 12x =144\\ x=12\] So, the amount of protein will be the same in Tori's drink if you increase the volume to 16 oz. So, Compared to Brady's, they have the same quantity of protein each, so they are in a 1:1 ratio

OpenStudy (kirbykirby):

Another way to see this... If you try 9/12 = 0.75 and 12/16 = 0.75 The amount of protein is found in the same concentration in both drinks.

OpenStudy (antonio_xx2):

that really helps a lot thanks

OpenStudy (kirbykirby):

:)